# Advances in the Shannon Capacity of Graphs

### v2

@@ Abstract
We derive exact values and new bounds for the Shannon capacity of two families of graphs: the $q$-Kneser graphs and the tadpole graphs. We also construct a countably infinite family of connected graphs whose Shannon capacity is not attained by the independence number of any finite strong power. Building on recent work of Schrijver, we establish sufficient conditions under which the Shannon capacity of a polynomial in graphs, formed via disjoint unions and strong products, equals the corresponding polynomial of the individual capacities, thereby reducing the evaluation of such capacities to that of their components. Finally, we prove an inequality relating the Shannon capacities of the strong product of graphs and their disjoint union, which yields streamlined proofs of several known bounds. In addition to contributing to the computation of the Shannon capacity of graphs, this paper is intended to serve as an accessible entry point to those wishing to work in this area.
Keywords. Shannon capacity of graphs, zero-error information theory, strong graph product; graph polynomials, independence number, Kneser graphs, tadpole graphs.
2020 Mathematics Subject Classification. 94A15 (primary), 05C35, 05C50, 05C69, 05C76 (secondary).
@@ 1 Introduction
The calculation of the Shannon capacity of a graph is notoriously difficult (see, e.g., [7, 8, 9, 10]), and only the capacity of a some families of graphs are known (e.g., Kneser graphs [11]). The computational complexity of the Shannon capacity has extended the research towards finding computable bounds on the capacity and exploring its properties [21, 20, 13, 14, 15, 16, 17, 11, 18, 19].
While exploring properties of the Shannon capacity, it is worth keeping in mind the more general concept of the asymptotic spectrum of graphs, introduced by Zuiddam [22, 23, 21, 20], which delineates a space of graph parameters that remain invariant under graph isomorphisms. This space is characterized by the following unique properties: additivity under disjoint union of graphs, multiplicativity under strong product of graphs, normalization for a simple graph with a single vertex, and monotonicity under graph complement homomorphisms. Building upon Strassen’s theory of the asymptotic spectra [24], a novel dual characterization of the Shannon capacity of a graph is derived in [22], expressed as the minimum over the elements of its asymptotic spectrum. By confirming that various graph invariants, including the Lovász $\vartheta$-function [11] and the fractional Haemers’ bound [14], are elements of the asymptotic spectrum of a graph (spectral points), it can be deduced that these elements indeed serve as upper bounds on the Shannon capacity of a graph. For further exploration, the comprehensive paper by Wigderson and Zuiddam [23] provides a survey on Strassen’s theory of the asymptotic spectra and its application areas, including the Shannon capacity of graphs.
@@
This paper studies several research directions regarding the Shannon capacity of graphs, and it is structured as follows.
@@
  • Section 6 derives the Shannon capacity of the $q$-Kneser graphs, using a generalized result of Erdős-Ko-Rado [29], and a result on the spectrum of the $q$-Kneser graphs [30]. This broadens the class of graphs for which the Shannon capacity is explicitly known.
  • Section 7 introduces a new inequality relating the Shannon capacity of the strong power of graphs to that of their disjoint union, and identifies several conditions under which equality holds. As an illustration of the applicability of this inequality, this section also presents an alternative shorter proof of a lower bound for the disjoint union of a graph with its complement, originally obtained by N. Alon [7], along with analogous bounds on the Lovász $\vartheta$-function of graphs.
  • Section 8 suggests some directions for further research that are related to the findings in this paper.
@@ 2.1.1 Terminology
Let $\mathsf{G}=(\operatorname{\mathsf{V}},\operatorname{\mathsf{E}})$ be a graph.
  • $\operatorname{\mathsf{V}}=\operatorname{\mathsf{V}}(\mathsf{G})$ is the vertex set of $\mathsf{G}$, and $\operatorname{\mathsf{E}}=\operatorname{\mathsf{E}}(\mathsf{G})$ is the edge set of $\mathsf{G}$.
@@
  • Vertices $i,j\in\operatorname{\mathsf{V}}(\mathsf{G})$ are adjacent if they are the endpoints of an edge in $\mathsf{G}$; it is denoted by $\{i,j\}\in\operatorname{\mathsf{E}}(\mathsf{G})$ or $i\sim j$.
@@ 2.1.2 Graph operations
Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. The strong product $\mathsf{G}\boxtimes\mathsf{H}$ is a graph whose vertices set is $\operatorname{\mathsf{V}}(\mathsf{G})\times\operatorname{\mathsf{V}}(\mathsf{H})$, and two distinct vertices $(g_{1},h_{1}),(g_{2},h_{2})$ are adjacent if one of the following three conditions is satisfied:
@@
Throughout this paper, we also use disjoint union of graphs as the addition operation on graphs.
@@ 2.1.3 Basic graph invariants under isomorphism
The independence number of a graph $\mathsf{G}$, denoted $\alpha(\mathsf{G})$ is the order of a largest independent set in $\mathsf{G}$, i.e.
@@
The clique number of a graph $\mathsf{G}$, denoted $\operatorname{\omega}(\mathsf{G})$ is the order of a largest clique in $\mathsf{G}$, i.e.
@@
The chromatic number of $\mathsf{G}$, denoted $\operatorname{\chi}(\mathsf{G})$, is the smallest cardinality of an independent set partition of $\mathsf{G}$.
@@
The clique-cover number of $\mathsf{G}$, denoted $\operatorname{\sigma}(\mathsf{G})$, is the smallest number of cliques needed to cover all the vertices of $\mathsf{G}$. Hence, $\operatorname{\sigma}(\mathsf{G})=\operatorname{\chi}(\overline{\mathsf{G}})$.
We next provide required properties of the independence number of a graph.
@@
Let $\mathsf{G}$ be a simple graph, and let $\mathsf{H}_{1}$ and $\mathsf{H}_{2}$ be induced and spanning subgraphs of $\mathsf{G}$, respectively. Then
@@ 2.1.4 Fractional invariants of graphs
Four basic fractional invariants of graphs are next defined by using linear programming.
@@
For a simple graph $\mathsf{G}$,
@@
  • The fractional chromatic number of $\mathsf{G}$ is
@@
[37] Let $\mathsf{G}$ be a simple graph. Then
@@
Some properties of the fractional independence number are next presented.
@@
[32] Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. Then
@@ 2.1.5 Graph spectrum
Let $\mathsf{G}$ be a simple undirected graph on $n$ vertices. The adjacency matrix of $\mathsf{G}$, denoted by $\operatorname{\mathbf{A}}=\operatorname{\mathbf{A}}(\mathsf{G})$, is an $n\times n$ symmetric matrix $\operatorname{\mathbf{A}}=(\mathrm{A}_{i,j})$ where $\mathrm{A}_{i,j}=1$ if $\{i,j\}\in\operatorname{\mathsf{E}}(\mathsf{G})$, and $\mathrm{A}_{i,j}=0$ otherwise (so, the entries in the principal diagonal of $\operatorname{\mathbf{A}}$ are zeros).
@@ 2.1.6 Some structured families of graphs
A graph $\mathsf{G}$ is said to be vertex-transitive if, for every two vertices $i,j\in\operatorname{\mathsf{V}}(\mathsf{G})$, there is an automorphism $f\colon\operatorname{\mathsf{V}}(\mathsf{G})\to\operatorname{\mathsf{V}}(\mathsf{G})$ such that $f(i)=j$.
@@
A graph $\mathsf{G}$ is edge-transitive if, for every two edges $e_{1},e_{2}\in\operatorname{\mathsf{E}}(\mathsf{G})$, there is an automorphism $f\colon\operatorname{\mathsf{V}}(\mathsf{G})\to\operatorname{\mathsf{V}}(\mathsf{G})$ that maps the endpoints of $e_{1}$ to the endpoints of $e_{2}$.
@@
A graph $\mathsf{G}$ is strongly regular with parameters $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ if it satisfies
@@
Let $q$ be a prime power, i.e. $q=p^{n}$ where $p$ is prime and $n\in\mathbb{N}$, with $q\equiv 1\text{ mod 4}$. The Paley graph of order $q$, denoted $P(q)$, is defined as
@@
[33] Let $q\equiv 1\text{ mod 4}$ be a prime power, and let $\mathsf{G}=P(q)$ be a Paley graph. Then,
@@ 2.2 The Shannon capacity of graphs
The concept of the Shannon capacity of a graph $\mathsf{G}$ was introduced in [1] to consider the largest information rate that can be achieved with zero-error communication. A discrete memoryless channel (DMC) with finite input and output alphabets is specified by a finite input set $\mathcal{X}$, a finite output set $\mathcal{Y}$, and a non-empty fan-out set $\mathcal{S}_{x}\triangleq\bigl\{y\in\mathcal{Y}:\,p_{Y|X}(y|x)>0\bigr\}$ for each $x\in\mathcal{X}$. In each channel use, the sender transmits an input $x\in\mathcal{X}$, and the receiver observes an arbitrary output in the corresponding fan-out set $\mathcal{S}_{x}$. It is possible to represent the channel by a confusion graph $\mathsf{G}$, which is defined as follows:
@@
for some index $i\in[k]$. Thus, it is possible to represent the super-channel by the $k$-th confusion graph that is defined as the $k$-fold strong power of $\mathsf{G}$. Using the $k$-th confusion graph, the largest amount of information that can be sent by $k$ uses of the channel with error-free communication is given by the independence number of the $k$-th confusion graph, i.e., it is equal to $\alpha(\mathsf{G}^{k})$. The maximum information rate per symbol that is achieved by $k$ uses of the channel, and with zero-error communication, is equal to
@@
We next provide several properties of the Shannon capacity, which are used throughout this paper.
@@
[23, 28] Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. Then
@@
Let $\mathsf{G}$ be a simple graph, and let $\mathsf{H}_{1}$ and $\mathsf{H}_{2}$ be induced and spanning subgraphs of $\mathsf{G}$, respectively. Then
@@
This holds by the definitions of induced and spanning subgraphs, and the definition in (27). ∎
@@ 2.3 The Lovász $\vartheta$-function of graphs
where the minimum is taken over
  • all orthonormal representations $\{\mathbf{u}_{i}:i\in\operatorname{\mathsf{V}}(\mathsf{G})\}$ of $\mathsf{G}$.
  • all unit vectors $\mathbf{c}$.
The unit vector $\mathbf{c}$ that achieves the minimum is called the handle of the orthonormal representation.
@@
Adding the inequality constraints $B_{i,j}\geq 0$ for all $i,j\in[n]$ to (40) yields the Schrijver $\vartheta$-function of $\mathsf{G}$, denoted $\vartheta^{\prime}(\mathsf{G})$, which therefore yields
@@
We next provide an alternative representation of the Lovász $\vartheta$-function of a graph [11].
@@
[11] Let $(\mathbf{v}_{1},\cdots,\mathbf{v}_{n})$ range over all orthonormal representations of $\overline{\mathsf{G}}$ and $\mathbf{d}$ over all unit vectors. Then,
@@
[11] Let $\mathsf{G}_{1}$ and $\mathsf{G}_{2}$ be simple graphs. Then
@@
[17] Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. Then
@@
Let $\mathsf{G}$ be a simple graph, and let $\mathsf{H}_{1}$ and $\mathsf{H}_{2}$ be induced and spanning subgraphs of $\mathsf{G}$, respectively. Then
@@
[11, 36] Let $\mathsf{G}$ be a simple graph of order $n$, then
@@
[11] Let $\mathsf{G}$ be a $d$-regular graph of order $n$, and let $\lambda_{n}$ be its smallest eigenvalue. Then

### v3

@@ Abstract
We derive exact values and new bounds for the Shannon capacity of two families of graphs: the $q$-Kneser graphs and the tadpole graphs. We also construct a countably infinite family of connected graphs whose Shannon capacity is not attained by the independence number of any finite strong power. Building on recent work of Schrijver, we establish sufficient conditions under which the Shannon capacity of a polynomial in graphs, formed via disjoint unions and strong products, equals the corresponding polynomial of the individual capacities, thereby reducing the evaluation of such capacities to that of their components. Finally, we prove an inequality relating the Shannon capacities of the strong product of graphs and their disjoint union, which yields alternative proofs of several known bounds as well as new tightness conditions. In addition to contributing to the computation of the Shannon capacity of graphs, this paper is intended to serve as an accessible entry point to those wishing to work in this area.
Keywords. Shannon capacity of graphs, zero-error information theory, strong graph product, independence number, Kneser graphs, tadpole graphs.
2020 Mathematics Subject Classification. 05C35, 05C50, 05C69, 05C76, 94A15.
@@ 1 Introduction
The calculation of the Shannon capacity of a graph is notoriously difficult (see, e.g., [7, 8, 9, 10]), and only the capacity of a some families of graphs are known (e.g., Kneser graphs [11]). The computational complexity of the Shannon capacity has extended the research towards finding computable bounds on the capacity and exploring its properties [7, 21, 20, 13, 14, 15, 16, 17, 11, 18, 19].
While exploring properties of the Shannon capacity, it is worth keeping in mind the more general concept of the asymptotic spectrum of graphs, introduced by Zuiddam [22, 23, 21, 20], which delineates a space of graph parameters that remain invariant under graph isomorphisms. This space is characterized by the following unique properties: additivity under disjoint union of graphs, multiplicativity under strong product of graphs, normalization for a simple graph with a single vertex, and monotonicity under graph complement homomorphisms. Building upon Strassen’s theory of asymptotic spectra [24], a novel dual characterization of the Shannon capacity of a graph is derived in [22], expressed as the minimum over the elements of its asymptotic spectrum. By confirming that various graph invariants, including the Lovász $\vartheta$-function [11] and the fractional Haemers’ bound [14], are elements of the asymptotic spectrum of a graph (spectral points), it can be deduced that these elements indeed serve as upper bounds on the Shannon capacity of a graph. For further exploration, the comprehensive paper by Wigderson and Zuiddam [23] provides a survey on Strassen’s theory of the asymptotic spectra and its application areas, including the Shannon capacity of graphs.
@@
This paper addresses several research directions in the study of the Shannon capacity of graphs, and it is organized as follows.
@@
  • Section 6 determines the Shannon capacity of the $q$-Kneser graphs in a manner analogous to the Lovász calculation of the Shannon capacity of the classical Kneser graphs [11]. The derivation relies on two known results, namely a generalized Erdős–Ko–Rado theorem for finite vector spaces and characterizations of the spectrum of the $q$-Kneser graphs. This broadens the class of graphs for which the exact Shannon capacity is explicitly determined.
  • Section 7 introduces a new inequality that relates the Shannon capacity of the strong product of graphs to that of their disjoint union, and it identifies several conditions under which equality is attained. As an illustration of the applicability of this inequality, the section also derives an alternative proof of a lower bound on the Shannon capacity of the disjoint union of a graph and its complement, originally due to Alon [7], together with new sufficient conditions for its attainability, as well as analogous bounds for the Lovász $\vartheta$-function.
  • Section 8 outlines directions for future research that are largely related to the results of this paper.
@@ 2.1.1 Terminology
Let $\mathsf{G}=(\operatorname{\mathsf{V}},\operatorname{\mathsf{E}})$ be a graph, where $\operatorname{\mathsf{V}}=\operatorname{\mathsf{V}}(\mathsf{G})$ is the vertex set of $\mathsf{G}$, and $\operatorname{\mathsf{E}}=\operatorname{\mathsf{E}}(\mathsf{G})$ is the edge set of $\mathsf{G}$.
@@
  • Vertices $i,j\in\operatorname{\mathsf{V}}(\mathsf{G})$ are adjacent if they are the endpoints of an edge in $\mathsf{G}$, which is denoted by $\{i,j\}\in\operatorname{\mathsf{E}}(\mathsf{G})$ or $i\sim j$.
@@ 2.1.2 Graph operations
Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. The strong product $\mathsf{G}\boxtimes\mathsf{H}$ is a graph whose vertices set is $\operatorname{\mathsf{V}}(\mathsf{G})\times\operatorname{\mathsf{V}}(\mathsf{H})$, and two distinct vertices $(g_{1},h_{1}),(g_{2},h_{2})$ are adjacent if, in each coordinate, they are either equal or adjacent. This means that one of the following three conditions is needed to be satisfied:
@@
Throughout this paper, we also use the disjoint union of graphs as the addition operation on graphs.
@@ 2.1.3 Basic graph invariants under isomorphism
The independence number of a graph $\mathsf{G}$, denoted by $\alpha(\mathsf{G})$, is the order of a largest independent set in $\mathsf{G}$, i.e.
@@
The clique number of a graph $\mathsf{G}$, denoted by $\operatorname{\omega}(\mathsf{G})$, is the order of a largest clique in $\mathsf{G}$, i.e.
@@
The chromatic number of a graph $\mathsf{G}$, denoted by $\operatorname{\chi}(\mathsf{G})$, is the smallest cardinality of a partition of the vertex set of $\mathsf{G}$ into independent sets. Equivalently, it is the minimum number of colors needed to color the vertices of $\mathsf{G}$ such that the endpoints of every edge receive distinct colors.
@@
The clique-cover number of $\mathsf{G}$, denoted by $\operatorname{\sigma}(\mathsf{G})$, is the smallest number of cliques needed to cover all the vertices of $\mathsf{G}$. Hence, $\operatorname{\sigma}(\mathsf{G})=\operatorname{\chi}(\overline{\mathsf{G}})$.
Next, we provide required properties of the independence number of a graph.
@@
Let $\mathsf{G}$ be a simple graph, and let $\mathsf{H}_{1}$ and $\mathsf{H}_{2}$ be induced and spanning subgraphs of $\mathsf{G}$, respectively. Then,
@@ 2.1.4 Fractional invariants of graphs
Next, four basic fractional invariants of graphs are defined by using linear programming.
@@
For a finite, undirected, and simple graph $\mathsf{G}$, fractional invariants of graphs are defined as follows.
@@
    Namely, $\alpha_{\mathrm{f}}(\mathsf{G})$ is the supremum of the total weight assigned to the vertices of $\mathsf{G}$, where nonnegative real weights are assigned so that for every clique of $\mathsf{G}$, the sum of the weights of the vertices in that clique is at most 1.
@@
    Namely, $\operatorname{\sigma}_{\mathrm{f}}(\mathsf{G})$ is the infimum of the total weight assigned to the cliques of $\mathsf{G}$, where nonnegative real weights are assigned so that for every vertex of $\mathsf{G}$, the sum of the weights of all cliques containing that vertex is at least 1.
@@
    Namely, $\operatorname{\omega}_{\mathrm{f}}(\mathsf{G})$ is the supremum of the total weight assigned to the vertices of $\mathsf{G}$, where nonnegative real weights are assigned so that for every independent set of $\mathsf{G}$, the sum of the weights of the vertices in that independent set is at most 1.
  • The fractional chromatic number of $\mathsf{G}$, denoted by $\operatorname{\chi}_{\mathrm{f}}(\mathsf{G})$, is given by
@@
    Namely, $\operatorname{\chi}_{\mathrm{f}}(\mathsf{G})$ is the infimum of the total weight assigned to the independent sets of $\mathsf{G}$, where nonnegative real weights are assigned so that for every vertex of $\mathsf{G}$, the sum of the weights of all independent sets containing that vertex is at least 1.
@@
[37] Let $\mathsf{G}$ be a simple graph. Then,
@@
Some properties of the fractional independence number are presented next.
@@
[32] Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. Then,
@@ 2.1.5 Graph spectrum
Let $\mathsf{G}$ be a simple undirected graph on $n$ vertices. The adjacency matrix of $\mathsf{G}$, denoted by $\operatorname{\mathbf{A}}=\operatorname{\mathbf{A}}(\mathsf{G})$, is an $n\times n$ symmetric matrix $\operatorname{\mathbf{A}}=(\mathrm{A}_{i,j})$ where $\mathrm{A}_{i,j}=1$ if $\{i,j\}\in\operatorname{\mathsf{E}}(\mathsf{G})$, and $\mathrm{A}_{i,j}=0$ otherwise (so the entries in the principal diagonal of $\operatorname{\mathbf{A}}$ are zeros).
@@ 2.1.6 Some structured families of graphs
A graph $\mathsf{G}$ is said to be vertex-transitive if for every two vertices $i,j\in\operatorname{\mathsf{V}}(\mathsf{G})$, there is an automorphism $f\colon\operatorname{\mathsf{V}}(\mathsf{G})\to\operatorname{\mathsf{V}}(\mathsf{G})$ such that $f(i)=j$.
@@
A graph $\mathsf{G}$ is edge-transitive if for every two edges $e_{1},e_{2}\in\operatorname{\mathsf{E}}(\mathsf{G})$, there is an automorphism $f\colon\operatorname{\mathsf{V}}(\mathsf{G})\to\operatorname{\mathsf{V}}(\mathsf{G})$ that maps the endpoints of $e_{1}$ to the endpoints of $e_{2}$.
@@
A graph $\mathsf{G}$ is strongly regular with parameters $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ if it satisfies the following:
@@
Let $q=p^{n}$ be a prime power with $p$ prime, $n\in\mathbb{N}$, and $q\equiv 1\text{ mod 4}$. The Paley graph of order $q$, denoted by $P(q)$, is defined as follows:
@@
[33] Let $q\equiv 1\text{ mod 4}$ be a prime power, and let $P(q)$ be a Paley graph. Then, the following hold:
@@ 2.2 The Shannon capacity of graphs
The Shannon capacity of a graph $\mathsf{G}$ was introduced in [1] as a measure of the largest effective alphabet size permitting zero-error communication over a given noisy channel. A discrete memoryless channel (DMC) with finite input and output alphabets is specified by a finite input set $\mathcal{X}$, a finite output set $\mathcal{Y}$, and a non-empty fan-out set $\mathcal{S}_{x}\triangleq\bigl\{y\in\mathcal{Y}:\,p_{Y|X}(y|x)>0\bigr\}$ for each $x\in\mathcal{X}$. In each channel use, the sender transmits an input $x\in\mathcal{X}$, and the receiver observes an arbitrary output in the corresponding fan-out set $\mathcal{S}_{x}$. The channel can be represented by a confusion graph $\mathsf{G}$ that is defined as follows:
@@
for some index $i\in[k]$. Thus, it is possible to represent the super-channel by the $k$th confusion graph that is defined as the $k$-fold strong power of $\mathsf{G}$. Using the $k$th confusion graph, the largest amount of information that can be sent by $k$ uses of the channel with error-free communication is given by the independence number of the $k$th confusion graph, i.e., it is equal to $\alpha(\mathsf{G}^{k})$. The maximum information rate per symbol that is achieved by $k$ uses of the channel, and with zero-error communication, is equal to
@@
Next, we provide several properties of the Shannon capacity, which are used throughout this paper.
@@
[23, 28] Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. Then,
@@
Let $\mathsf{G}$ be a simple graph, and let $\mathsf{H}_{1}$ and $\mathsf{H}_{2}$ be induced and spanning subgraphs of $\mathsf{G}$, respectively. Then,
@@
This holds by the definitions of induced and spanning subgraphs, together with (27). ∎
@@ 2.3 The Lovász $\vartheta$-function of graphs
where the minimum is taken over all orthonormal representations $\{\mathbf{u}_{i}:i\in\operatorname{\mathsf{V}}(\mathsf{G})\}$ of $\mathsf{G}$ and all unit vectors $\mathbf{c}$. The unit vector $\mathbf{c}$ attaining the minimum is called the handle of the orthonormal representation.
@@
Adding the inequality constraints $B_{i,j}\geq 0$ for all $i,j\in[n]$ to (40) yields the Schrijver $\vartheta$-function of $\mathsf{G}$, denoted by $\vartheta^{\prime}(\mathsf{G})$, which therefore yields
@@
Next, we provide an alternative representation of the Lovász $\vartheta$-function of a graph [11].
@@
[11] Let $(\mathbf{v}_{1},\cdots,\mathbf{v}_{n})$ range over all orthonormal representations of $\overline{\mathsf{G}}$, and let $\mathbf{d}$ range over all unit vectors. Then,
@@
[11] Let $\mathsf{G}_{1}$ and $\mathsf{G}_{2}$ be simple graphs. Then,
@@
[17] Let $\mathsf{G}$ and $\mathsf{H}$ be simple graphs. Then,
@@
Let $\mathsf{G}$ be a simple graph, and let $\mathsf{H}_{1}$ and $\mathsf{H}_{2}$ be induced and spanning subgraphs of $\mathsf{G}$, respectively. Then,
@@
[11, 36] Let $\mathsf{G}$ be a simple graph on $n$ vertices. Then,
@@
[11] Let $\mathsf{G}$ be a $d$-regular graph of order $n$, and let $\lambda_{n}$ be its smallest eigenvalue. Then,